# A deep backward regression-based scheme for high-dimensional nonlinear partial differential equations

### 16 Mar 2026 v1

@@ 3.3 Convergence analysis
$\max\limits_{0\leq i\leq N-1}\mathbb{E}\left[\left|Y_{t_{i}}-\mathcal{Y}_{i}\right|^{2}\right]+\mathbb{E}\left[\sum\limits_{i=0}^{N-1}\int_{t_{i}}^{t_{i+1}}|Z_{s}-\mathcal{Z}_{i}|^{2}ds\right]\leq Ch,$
which completes the proof.
@@ 4 Extension to variational inequalities
In this section, we extend the DBR scheme to variational inequalities, referring to the resulting method as the RDBR scheme. Then we study the convergence of the RDBR scheme for the variational inequalities.
@@ 4.1 RDBR Scheme
$\begin{cases}\min\left(-\partial_{t}u-\mu\cdot D_{x}u-\frac{1}{2}\operatorname{Tr}\left(\sigma\sigma^{\top}D_{x}^{2}u\right)+f\left(\cdot,\cdot,u,\sigma^{\top}D_{x}u\right),u-g\right)=0,&\text{ on }[0,T)\times\mathbb{R}^{d},\\ u(T,x)=g(x),&\text{ on }R^{d}.\end{cases}$ (4.1)
This inequality arises, for instance, in optimal stopping problems and American option pricing. It is established that such a variational inequality is associated with a reflected backward stochastic differential equation (RBSDE) given by
$\begin{cases}X_{t}=x_{0}+\int_{0}^{t}\mu(s,X_{s})ds+\int_{0}^{t}\sigma(s,X_{s})dW_{s},\\ Y_{t}=g(X_{T})+\int_{t}^{T}f(s,X_{s},Y_{s},Z_{s})ds-\int_{t}^{T}Z_{s}dW_{s}+K_{T}-K_{t},\\ Y_{t}\geq g(X_{t}),\quad\forall t\in[0,T],\\ \int_{0}^{T}(Y_{t}-g(X_{t}))dK_{t}=0,\end{cases}$ (4.2)
where $K$ is an adapted non-decreasing process satisfying the Skorokhod condition (see [14]). The Euler time-discretization of the RFBSDE (4.2), at the mesh points $t_{i}$, is
$Y_{i}^{\pi}=Y_{i+1}^{\pi}+hf(t_{i},X_{i}^{\pi},Y_{i}^{\pi},Z_{i}^{\pi})-Z_{i}^{\pi}\Delta W_{i}+K_{i+1}^{\pi}-K_{i}^{\pi},$
which also reads as the conditional expectation formula
$\left\{\begin{aligned} \overline{Y}_{i}^{\pi}=&\mathbb{E}_{i}\left[Y^{\pi}_{i+1}+hf(t_{i},X_{i}^{\pi},\overline{Y}_{i}^{\pi},Z_{i}^{\pi})\right],\\ Y_{i}^{\pi}=&\max\left\{\overline{Y}^{\pi}_{i},g\left(X_{i}^{\pi}\right)\right\},\\ Z_{i}^{\pi}=&\mathbb{E}_{i}\left[Y_{i+1}^{\pi}\frac{\Delta W_{i}^{\top}}{h}\right].\end{aligned}\right$.
The backward process $Y_{\cdot}^{\pi,m}$ in (4.2) is simulated by Monte Carlo method in the following ways, for $i=N-1,N-2,\cdots,0$
$Y_{i}^{\pi,m}=Y_{i+1}^{\pi,m}+hf(t_{i},X_{i}^{\pi,m},Y_{i}^{\pi,m},Z_{i}^{\pi,m})-Z_{i}^{\pi,m}\Delta W_{i}^{m}+K_{i+1}^{\pi,m}-K_{i}^{\pi,m},$
which also expresses as the conditional expectation form
$\left\{\begin{aligned} \overline{Y}_{i}^{\pi,m}=&\mathbb{E}_{i}\Big[Y_{i+1}^{\pi,m}+hf(t_{i},X_{i}^{\pi,m},\overline{Y}_{i}^{\pi,m},Z_{i}^{\pi,m})\Big],\\ Y_{i}^{\pi,m}=&\max\left\{\overline{Y}^{\pi,m}_{i},g\left(X_{i}^{\pi,m}\right)\right\},\\ Z_{i}^{\pi,m}=&\mathbb{E}_{i}\big[Y_{i+1}^{\pi,m}\frac{(\Delta W_{i}^{m})^{\top}}{h}\big].\end{aligned}\right$. (4.3)
Thus, we derive the fully discrete approximation $(\widehat{\mathcal{Y}}_{i},\widehat{\mathcal{Z}}_{i})$ for the RDBR scheme for the solutions $(Y_{t_{i}},Z_{t_{i}})$ of the variational inequality (4.1) as follows, for $i=N,N-1,\cdots,1,0$, $m=1,2,\cdots,1,M$
@@
    the terminal condition is $(\widehat{\mathcal{Y}}_{N},\widehat{\mathcal{Z}}_{N})=(\Phi(X_{N}^{\pi,m}),\sigma^{\top}(t_{N},X^{\pi,m}_{N})D_{x}u(t_{N},X^{\pi,m}_{N}))$,
@@
    $\left\{\begin{aligned} \widehat{\mathcal{Y}}_{i}=&\max\left\{\widehat{\mathcal{NN}}_{y,i,1+d,n}^{\varrho,L,\mathbf{n}}(p^{m}_{i},\theta^{*}_{y,i}),g\left(X_{i}^{\pi,m}\right)\right\},\\ \widehat{\mathcal{Z}}_{i}=&\widehat{\mathcal{NN}}_{z,i,1+d,n\times d}^{\varrho,L,\mathbf{n}}(p^{m}_{i},\theta^{*}_{z,i}).\end{aligned}\right$. (4.4)
    and compute (by SGD) the minimizer of the quadratic loss function, for $r_{j}\in\mathbb{N}^{+}$ large enough
    $\left\{\begin{aligned} \widehat{F}_{y,i}^{r_{j}}(\theta_{y,i}^{r_{j}})=&\frac{1}{M}\sum_{m=1}^{M}\left|S_{y,i}\left(p_{i}^{m},X_{i+1}^{\pi,m}\right)-\widehat{\mathcal{NN}}_{y,i,1+d,n}^{\varrho,L,\mathbf{n}}(p_{i}^{m},\theta_{y,i}^{r_{j}})\right|^{2},\\ \widehat{F}_{z,i}^{r_{j}}(\theta_{z,i}^{r_{j}})=&\frac{1}{M}\sum_{m=1}^{M}\left|S_{z,i}\left(p_{i}^{m},\Delta W_{i}^{m},X_{i+1}^{\pi,m}\right)-\widehat{\mathcal{NN}}_{z,i,1+d,n\times d}^{\varrho,L,\mathbf{n}}(p_{i}^{m},\theta_{z,i}^{r_{j}})\right|^{2},\end{aligned}\right$.
    where $\theta^{*}_{y,i}\in\arg\min\limits_{\theta_{y,i}^{r_{j}}}\widehat{F}_{y,i}^{r_{j}}(\theta_{y,i}^{r_{j}})$, $\theta^{*}_{z,i}\in\arg\min\limits_{\theta_{z,i}^{r_{j}}}\widehat{F}_{z,i}^{r_{j}}(\theta_{z,i}^{r_{j}})$.
@@ 4.2 Convergence analysis
In this section, we analyze the convergence of the RDBR schemes (4.4) for the variational inequality (4.1) related to the solution $(Y,Z)$ of the RFBSDE (4.2).
@@
Now, we investigate the errors of the scheme (4.4) and define the auxiliary quantities as below, for $i=N-1,N-2,\cdots,0$
@@
$\left\{\begin{aligned} \overline{\mathcal{U}}_{t_{i}}^{m}&=\mathbb{E}_{i}\left[\widehat{\mathcal{NN}}_{y,i+1,1+d,n}^{\varrho,L,\mathbf{n}}(p_{i+1}^{m},\theta^{*}_{y,i+1})+hf\left(t_{i},X_{i}^{\pi,m},\overline{\mathcal{U}}_{t_{i}}^{m},\widehat{\mathcal{V}}_{t_{i}}^{m}\right)\right],\\ \widehat{\mathcal{U}}_{t_{i}}^{m}&=\max\left\{\overline{\mathcal{U}}_{t_{i}}^{m},g\left(X_{i}^{\pi,m}\right)\right\},\\ \widehat{\mathcal{V}}_{t_{i}}^{m}&=\mathbb{E}_{i}\left[\widehat{\mathcal{NN}}_{y,i+1,1+d,n}^{\varrho,L,\mathbf{n}}(p_{i+1}^{m},\theta^{*}_{y,i+1})\frac{(\Delta W_{i}^{m})^{\top}}{h}\right].\end{aligned}\right$. (4.5)
The result is obtained under one of the following additional assumptions:
Assumption 4.1
$g$ is $C^{1}$, and $g,D_{x}g$ are Lipschitz.
Assumption 4.2
$\sigma$ is $C^{1}$, with $\sigma,D_{x}\sigma$ both Lipschitz, and $g$ is $C^{2}$, with $g,D_{x}g,D^{2}_{x}g$ all Lipschitz.
Theorem 4.3
Let the assumptions (i)-(iv) hold. Let $(Y_{t_{i}},Z_{t_{i}})$ and $(\widehat{\mathcal{Y}}_{i},\widehat{\mathcal{Z}}_{i})$ be solutions of the RFBSDE (4.2) and solutions of the RDBR method (4.4) respectively. Then, we have, for small enough $h$
$\max\limits_{0\leq i\leq N-1}\mathbb{E}\left[\left|Y_{t_{i}}-\widehat{\mathcal{Y}}_{i}\right|^{2}\right]+\mathbb{E}\left[\sum\limits_{i=0}^{N-1}\int_{t_{i}}^{t_{i+1}}|Z_{s}-\widehat{\mathcal{Z}}_{i}|^{2}ds\right]$
$\leq C\varepsilon(\pi)+C\sum\limits_{i=0}^{N-1}\mathbb{E}\left[N\left|\widehat{R}_{i+1,2}^{y}\right|^{2}+\left|\widehat{R}_{i,2}^{z}\right|^{2}+N\left|\widehat{R}_{i+1,1}^{y}\right|^{2}+\left|\widehat{R}_{i,1}^{z}\right|^{2}\right],$
with $\varepsilon(h)=O(h^{1/2})$ under Assumption 4.1, and $\varepsilon(h)=O(h)$ under Assumption 4.2;
$\widehat{R}_{i,1}^{y}$ $=\mathbb{E}_{i}\Big[Y_{i+1}^{\pi,m}+hf(t_{i},X_{i}^{\pi,m},\overline{Y}_{i}^{\pi,m},Z_{i}^{\pi,m})\Big]-\frac{1}{K_{i}}\sum\limits_{k=1}^{K_{i}}Y_{i+1}^{\pi,m,k}-hf(t_{i},X_{i}^{\pi,m},\overline{Y}_{i}^{\pi,m},Z_{i}^{\pi,m}),$
@@
$\widehat{R}_{i,1}^{z}$ $=\mathbb{E}_{i}\left[Y_{i+1}^{\pi,m}\frac{(\Delta W_{i}^{m})^{\top}}{h}\right]-\frac{1}{K_{i}}\sum\limits_{k=1}^{K_{i}}\left[Y_{i+1}^{\pi,m,k}\frac{\left(\Delta W_{i}^{m,k}\right)^{\top}}{h}\right],$
$\widehat{R}_{i,2}^{y}$ $=\frac{1}{K_{i}}\sum\limits_{k=1}^{K_{i}}Y_{i+1}^{\pi,m,k}+hf(t_{i},X_{i}^{\pi,m},\overline{Y}_{i}^{\pi,m},Z_{i}^{\pi,m})-\widehat{\mathcal{NN}}_{y,i,1+d,n}^{\varrho,L,\mathbf{n}}(p_{i}^{m},\theta_{y,i}^{*}),$
$\widehat{R}_{i,2}^{z}$ $=\frac{1}{K_{i}}\sum\limits_{k=1}^{K_{i}}\left[Y_{i+1}^{\pi,m,k}\frac{\left(\Delta W_{i}^{m,k}\right)^{\top}}{h}\right]-\widehat{\mathcal{NN}}_{z,i,1+d,n\times d}^{\varrho,L,\mathbf{n}}(p_{i}^{m},\theta_{z,i}^{*}).$
Proof. From [3], we have
$\begin{cases}\max\limits_{0\leq i\leq N-1}\mathbb{E}\left[\left|Y_{t_{i}}-Y_{i}^{\pi,m}\right|^{2}\right]&=\varepsilon(h),\\ \mathbb{E}\left[\sum\limits_{i=0}^{N-1}\int_{t_{i}}^{t_{i+1}}|Z_{t}-Z^{\pi,m}_{i}|^{2}dt\right]&=O(h^{\frac{1}{2}}),\end{cases}$ (4.6)
with $\varepsilon(h)=O(h^{1/2})$ under Assumption 4.1, and $\varepsilon(h)=O(h)$ under Assumption 4.2. From (4.3) and (4.5), we have, for $i\in\{0,1,\ldots,N-1\}$
@@
$\overline{Y}_{i}^{\pi,m}-\overline{\mathcal{U}}^{m}_{t_{i}}=$ $\mathbb{E}_{i}\left[Y_{i+1}^{\pi,m}-\widehat{\mathcal{NN}}_{y,i+1,1+d,n}^{\varrho,L,\mathbf{n}}(p_{i+1}^{m},\theta^{*}_{y,i+1})\right]$ (4.7)
$+h\left(f(t_{i},X_{i}^{\pi,m},\overline{Y}_{i}^{\pi,m},Z_{i}^{\pi,m})-f\left(t_{i},X_{i}^{\pi,m},\overline{\mathcal{U}}_{t_{i}}^{m},\widehat{\mathcal{V}}_{t_{i}}^{m}\right)\right).$
Proceeding similarly as the Step 1 in the proof pf Theorem 3.1, we obtain by Young inequality, Cauchy-Schwarz inequality, and the Lipschitz condition with respect to $f$
@@
$\mathbb{E}\left[\left|\overline{Y}_{i}^{\pi,m}-\overline{\mathcal{U}}^{m}_{t_{i}}\right|^{2}\right]$ $\leq(1+\gamma h)\mathbb{E}\left[\left|\mathbb{E}_{i}\left[Y_{i+1}^{\pi,m}-\widehat{\mathcal{NN}}_{y,i+1,1+d,n}^{\varrho,L,\mathbf{n}}(p_{i+1}^{m},\theta^{*}_{y,i+1})\right]\right|^{2}\right]$

### 21 May 2026 v3

@@ 3.3 Convergence analysis
$\inf_{\mathcal{NN}_{y,i}\in\mathcal{G}_{\mathcal{K}_{i},1+d,n}^{\zeta_{\kappa},L,\mathbf{n}}}\mathbb{E}\left[\left|y_{i}(X_{i}^{\pi,m})-\mathcal{NN}_{y,i,1+d,n}^{\varrho,L,\mathbf{n}}(p_{i}^{m},\theta_{y,i})\right|^{2}\right]$ (77)
$\leq\mathbb{E}\left[\left|y_{i}(X_{i}^{\pi,m})-\mathcal{N}_{y,i}^{*}(X_{i}^{\pi,m})\right|^{2}\mathbf{1}_{\{|X_{i}^{\pi,m}|\leq R\}}\right]$
$\quad+\mathbb{E}\left[\left|y_{i}(X_{i}^{\pi,m})-\mathcal{N}_{y,i}^{*}(X_{i}^{\pi,m})\right|^{2}\mathbf{1}_{\{|X_{i}^{\pi,m}|>R\}}\right].$
From (75), the first term on the right-hand side of (77) is bounded by
$\mathbb{E}\left[\left|y_{i}(X_{i}^{\pi,m})-\mathcal{N}_{y,i}^{*}(X_{i}^{\pi,m})\right|^{2}\mathbf{1}_{\{|X_{i}^{\pi,m}|\leq R\}}\right]\leq CR^{2}\eta^{2}.$ (78)
For the tail term, by Hölder inequality,
$\mathbb{E}\left[\left|y_{i}(X_{i}^{\pi,m})-\mathcal{N}_{y,i}^{*}(X_{i}^{\pi,m})\right|^{2}\mathbf{1}_{\{|X_{i}^{\pi,m}|>R\}}\right]$ (79)
$\leq\left(\mathbb{E}\left[\left|y_{i}(X_{i}^{\pi,m})-\mathcal{N}_{y,i}^{*}(X_{i}^{\pi,m})\right|^{2q}\right]\right)^{\frac{1}{q}}\left(\mathbb{P}\left(|X_{i}^{\pi,m}|>R\right)\right)^{\frac{q-1}{q}}.$
Since both $y_{i}$ and $\mathcal{N}_{y,i}^{*}$ are Lipschitz uniformly in $i$ and $N$, and since
$\left|y_{i}(0)-\mathcal{N}_{y,i}^{*}(0)\right|\leq CR\eta,$
we have
$\left(\mathbb{E}\left[\left|y_{i}(X_{i}^{\pi,m})-\mathcal{N}_{y,i}^{*}(X_{i}^{\pi,m})\right|^{2q}\right]\right)^{\frac{1}{q}}\leq C\left(1+R^{2}\eta^{2}\right).$ (80)
Moreover, by Markov inequality and the standard moment estimate for the Euler scheme,
$\mathbb{P}\left(|X_{i}^{\pi,m}|>R\right)\leq\frac{\mathbb{E}\left[|X_{i}^{\pi,m}|^{2q}\right]}{R^{2q}}\leq\frac{C}{R^{2q}},$ (81)
where we used $X_{0}\in L^{2+\delta}$ and $2q=2+\delta$. Therefore,
$\left(\mathbb{P}\left(|X_{i}^{\pi,m}|>R\right)\right)^{\frac{q-1}{q}}\leq\frac{C}{R^{2(q-1)}}=\frac{C}{R^{\delta}}.$ (82)
Combining (79), (80), and (82), we get
$\mathbb{E}\left[\left|y_{i}(X_{i}^{\pi,m})-\mathcal{N}_{y,i}^{*}(X_{i}^{\pi,m})\right|^{2}\mathbf{1}_{\{|X_{i}^{\pi,m}|>R\}}\right]\leq C\left(1+R^{2}\eta^{2}\right)R^{-\delta}.$ (83)
Consequently, by (77), (78), and (83), for $R\geq 1$ and $R\eta\leq 1$, we obtain
$\varepsilon_{i}^{y}$ $\leq C\left(R^{2}\eta^{2}+R^{-\delta}\right).$ (84)
Next, we estimate the approximation error of the $Z$-component. Applying the same argument to the functions $\sqrt{h}\,z_{i}$ and $\sqrt{h}\,\mathcal{N}_{z,i}^{*}$, and using (76), we obtain
$h\varepsilon_{i}^{z}$ $=\inf_{\mathcal{NN}_{z,i}\in\mathcal{G}_{\sqrt{\frac{d}{h}}\mathcal{K}_{i},1+d,n\times d}^{\zeta_{\kappa},L,\mathbf{n}}}\mathbb{E}\left[h\left|z_{i}(X_{i}^{\pi,m})-\mathcal{NN}_{z,i,1+d,n\times d}^{\varrho,L,\mathbf{n}}(p_{i}^{m},\theta_{z,i})\right|^{2}\right]$ (85)
$\leq C\left(R^{2}\eta^{2}+R^{-\delta}\right).$
Set
$a_{R,\eta}=R^{2}\eta^{2}+R^{-\delta}.$ (86)
Then (84) and (85) imply
$\varepsilon_{i}^{y}\leq Ca_{R,\eta},\qquad h\varepsilon_{i}^{z}\leq Ca_{R,\eta}.$ (87)
Therefore,
$\sum_{i=0}^{N-1}\left(N\varepsilon_{i}^{y}+\varepsilon_{i}^{z}\right)\leq CN^{2}a_{R,\eta}.$ (88)
Let
$K_{\mathrm{MC}}:=\min_{0\leq i\leq N-1}K_{i}$
be the minimal number of inner Monte Carlo samples. To achieve the convergence rate $O(1/N)$ in Theorem 1, it is sufficient to choose $R,\eta,K_{\mathrm{MC}}$ such that
$N^{2}\left(R^{2}\eta^{2}+R^{-\delta}\right)=O\left(\frac{1}{N}\right),\qquad\frac{N^{2}}{K_{\mathrm{MC}}}=O\left(\frac{1}{N}\right).$ (89)
This is verified by taking
$R=O\left(N^{\frac{3}{\delta}}\right),\qquad\eta=O\left(N^{-\frac{3}{2}-\frac{3}{\delta}}\right),\qquad K_{\mathrm{MC}}=O(N^{3}).$ (90)
Indeed, under this choice,
$R^{2}\eta^{2}=O(N^{-3}),\qquad R^{-\delta}=O(N^{-3}),\qquad\frac{N^{2}}{K_{\mathrm{MC}}}=O(N^{-1}).$
From Lemma 2, if $d>1$, the GroupSort neural networks can be chosen with grouping size
$\kappa=O\left(\left\lceil 2\sqrt{d}\,N^{\frac{3}{2}+\frac{3}{\delta}}\right\rceil\right),$ (91)
depth $L+1=O(d^{2})$, and width
$\sum_{\ell=0}^{L-1}n_{\ell}=O\left(\left(2\sqrt{d}\,N^{\frac{3}{2}+\frac{3}{\delta}}\right)^{d^{2}-1}\right).$ (92)
If $d=1$, one can take
$\kappa=O\left(N^{\frac{3}{2}+\frac{3}{\delta}}\right),\qquad L+1=3,\qquad\sum_{\ell=0}^{L-1}n_{\ell}=O\left(N^{\frac{3}{2}+\frac{3}{\delta}}\right).$ (93)
Since $g$ is Lipschitz and
$\max_{0\leq i\leq N-1}\mathbb{E}\left[\sup_{s\in[t_{i},t_{i+1}]}|X_{s}-X_{i}^{\pi,m}|^{2}\right]\leq Ch$
see [31], we obtain
$\mathbb{E}\left[\left|g(X_{T})-g(X_{N}^{\pi,m})\right|^{2}\right]+\mathbb{E}\left[\sum_{i=0}^{N-1}\int_{t_{i}}^{t_{i+1}}|Z_{s}-\bar{Z}_{t_{i}}|^{2}ds\right]\leq Ch.$ (94)
Combining (84), (85), (89), and (94) with Theorem 1, we finally obtain
$\max_{0\leq i\leq N-1}\mathbb{E}\left[\left|Y_{t_{i}}-\mathcal{Y}_{i}\right|^{2}\right]+\mathbb{E}\left[\sum_{i=0}^{N-1}\int_{t_{i}}^{t_{i+1}}|Z_{s}-\mathcal{Z}_{i}|^{2}ds\right]\leq Ch,$
which completes the proof. ∎
@@ 4 Extension to variational inequalities
In this section, we extend the DBR scheme to variational inequalities, referring to the resulting method as the RDBR scheme. Then we study the convergence of the RDBR scheme for the variational inequalities. Since the obstacle constraint requires an order relation and a projection operator, we restrict this section to the scalar case $n=1$. Thus, in this section, $u,Y,g,\Phi$ are real-valued, $Z\in\mathbb{R}^{1\times d}$, and the reflecting process $\mathbf{K}$ is a real-valued adapted non-decreasing process. The maximum operator $\max\{\cdot,\cdot\}$ is understood in the usual scalar sense. We keep the boldface notation $\mathbf{K}$ in this section to distinguish the reflecting process from the Monte Carlo sample sizes $K_{i}$. We also assume the compatibility condition
$g(x)\geq\Phi(x),\qquad x\in\mathbb{R}^{d}.$
@@ 4.1 RDBR Scheme
$\begin{cases}\min\left(-\partial_{t}u-\mu\cdot D_{x}u-\frac{1}{2}\operatorname{Tr}\left(\sigma\sigma^{\top}D_{x}^{2}u\right)-f\left(\cdot,\cdot,u,\sigma^{\top}D_{x}u\right),u-\Phi\right)=0,&\text{on }[0,T)\times\mathbb{R}^{d},\\ u(T,x)=g(x),&\text{on }\mathbb{R}^{d}.\end{cases}$ (95)
This inequality arises, for instance, in optimal stopping problems and American option pricing. It is established that such a variational inequality is associated with a reflected forward backward stochastic differential equation (RFBSDE) given by
$\begin{cases}X_{t}=x_{0}+\displaystyle\int_{0}^{t}\mu(s,X_{s})ds+\displaystyle\int_{0}^{t}\sigma(s,X_{s})dW_{s},\\[2.84526pt] Y_{t}=g(X_{T})+\displaystyle\int_{t}^{T}f(s,X_{s},Y_{s},Z_{s})ds-\displaystyle\int_{t}^{T}Z_{s}dW_{s}+\mathbf{K}_{T}-\mathbf{K}_{t},\\[2.84526pt] Y_{t}\geq\Phi(X_{t}),\quad\forall t\in[0,T],\\[2.84526pt] \displaystyle\int_{0}^{T}\left(Y_{t}-\Phi(X_{t})\right)d\mathbf{K}_{t}=0,\end{cases}$ (96)
where $\Phi:\mathbb{R}^{d}\rightarrow\mathbb{R}$ denotes the reflecting barrier and $\mathbf{K}$ is a real-valued adapted non-decreasing process satisfying the Skorokhod condition (see [14]).
The Euler time-discretization of the RFBSDE (96), at the mesh points $t_{i}$, is
$Y_{i}^{\pi}=Y_{i+1}^{\pi}+hf(t_{i},X_{i}^{\pi},Y_{i}^{\pi},Z_{i}^{\pi})-Z_{i}^{\pi}\Delta W_{i}+\mathbf{K}_{i+1}^{\pi}-\mathbf{K}_{i}^{\pi}.$
Equivalently, it can be written in the following conditional expectation form:
$\left\{\begin{aligned} \overline{Y}_{i}^{\pi}=&\ \mathbb{E}_{i}\left[Y_{i+1}^{\pi}+hf(t_{i},X_{i}^{\pi},\overline{Y}_{i}^{\pi},Z_{i}^{\pi})\right],\\ Y_{i}^{\pi}=&\ \max\left\{\overline{Y}_{i}^{\pi},\Phi(X_{i}^{\pi})\right\},\\ Z_{i}^{\pi}=&\ \mathbb{E}_{i}\left[Y_{i+1}^{\pi}\frac{\Delta W_{i}^{\top}}{h}\right].\end{aligned}\right$. (97)
The backward process $Y_{\cdot}^{\pi,m}$ in (96) is simulated by the Monte Carlo method in the following way, for $i=N-1,N-2,\ldots,0$:
$Y_{i}^{\pi,m}=Y_{i+1}^{\pi,m}+hf(t_{i},X_{i}^{\pi,m},Y_{i}^{\pi,m},Z_{i}^{\pi,m})-Z_{i}^{\pi,m}\Delta W_{i}^{m}+\mathbf{K}_{i+1}^{\pi,m}-\mathbf{K}_{i}^{\pi,m}.$
This can also be expressed as the conditional expectation form
$\left\{\begin{aligned} \overline{Y}_{i}^{\pi,m}=&\ \mathbb{E}_{i}\left[Y_{i+1}^{\pi,m}+hf(t_{i},X_{i}^{\pi,m},\overline{Y}_{i}^{\pi,m},Z_{i}^{\pi,m})\right],\\ Y_{i}^{\pi,m}=&\ \max\left\{\overline{Y}_{i}^{\pi,m},\Phi(X_{i}^{\pi,m})\right\},\\ Z_{i}^{\pi,m}=&\ \mathbb{E}_{i}\left[Y_{i+1}^{\pi,m}\frac{(\Delta W_{i}^{m})^{\top}}{h}\right].\end{aligned}\right$. (98)
Thus, we define the fully discrete approximations $\{\widehat{\mathcal{Y}}_{i}\}_{i=0}^{N}$ and $\{\widehat{\mathcal{Z}}_{i}\}_{i=0}^{N-1}$ for the RDBR scheme associated with the solutions $(Y_{t_{i}},Z_{t_{i}})$ of the variational inequality (95), for $m=1,2,\ldots,M$, as follows:
@@
    the terminal condition is
    $\widehat{\mathcal{Y}}_{N}=g(X_{N}^{\pi,m}).$
    No terminal value $\widehat{\mathcal{Z}}_{N}$ is required by the backward scheme.
@@
    $\left\{\begin{aligned} \overline{\mathcal{Y}}_{i}=&\ \widehat{\mathcal{NN}}_{y,i,1+d,n}^{\varrho,L,\mathbf{n}}(p_{i}^{m},\theta_{y,i}^{*}),\\ \widehat{\mathcal{Y}}_{i}=&\ \max\left\{\overline{\mathcal{Y}}_{i},\Phi(X_{i}^{\pi,m})\right\},\\ \widehat{\mathcal{Z}}_{i}=&\ \widehat{\mathcal{NN}}_{z,i,1+d,n\times d}^{\varrho,L,\mathbf{n}}(p_{i}^{m},\theta_{z,i}^{*}).\end{aligned}\right$. (99)
For the training at time $t_{i}$, the next-step value used in the conditional expectation is the post-projection value. For $m=1,\ldots,M$ and $k=1,\ldots,K_{i}$, define
[…]
Then the empirical losses are given by
$\left\{\begin{aligned} \widehat{F}_{z,i}^{r_{j}}(\theta_{z,i}^{r_{j}})=&\ \frac{1}{M}\sum_{m=1}^{M}\left|\frac{1}{K_{i}}\sum_{k=1}^{K_{i}}\widehat{\mathcal{Y}}_{i+1}^{m,k}\frac{(\Delta W_{i}^{m,k})^{\top}}{h}-\widehat{\mathcal{NN}}_{z,i,1+d,n\times d}^{\varrho,L,\mathbf{n}}(p_{i}^{m},\theta_{z,i}^{r_{j}})\right|^{2},\\ \widehat{F}_{y,i}^{r_{j}}(\theta_{y,i}^{r_{j}})=&\ \frac{1}{M}\sum_{m=1}^{M}\Bigg|\frac{1}{K_{i}}\sum_{k=1}^{K_{i}}\widehat{\mathcal{Y}}_{i+1}^{m,k}+hf\Big(t_{i},X_{i}^{\pi,m},\widehat{\mathcal{NN}}_{y,i,1+d,n}^{\varrho,L,\mathbf{n}}(p_{i}^{m},\theta_{y,i}^{r_{j}}),\\ &\qquad\qquad\qquad\qquad\widehat{\mathcal{NN}}_{z,i,1+d,n\times d}^{\varrho,L,\mathbf{n}}(p_{i}^{m},\theta_{z,i}^{*})\Big)-\widehat{\mathcal{NN}}_{y,i,1+d,n}^{\varrho,L,\mathbf{n}}(p_{i}^{m},\theta_{y,i}^{r_{j}})\Bigg|^{2}.\end{aligned}\right$. (101)
Here
$\theta_{z,i}^{*}\in\arg\min_{\theta_{z,i}^{r_{j}}}\widehat{F}_{z,i}^{r_{j}}(\theta_{z,i}^{r_{j}}),\qquad\theta_{y,i}^{*}\in\arg\min_{\theta_{y,i}^{r_{j}}}\widehat{F}_{y,i}^{r_{j}}(\theta_{y,i}^{r_{j}}).$
As in Section 3, the losses $\widehat{F}_{y,i}^{r_{j}}$ and $\widehat{F}_{z,i}^{r_{j}}$ above are empirical losses used in the numerical implementation. The convergence analysis below is carried out at the population-loss level. That is, all approximation errors are measured with respect to the true expectation $\mathbb{E}[\cdot]$, and exact minimization of the corresponding population losses is assumed. The additional errors caused by empirical risk approximation and non-exact stochastic optimization are not included in Theorem 4.
@@ 4.2 Convergence analysis
In this section, we analyze the convergence of the RDBR scheme (99) for the variational inequality (95) related to the solution $(Y,Z)$ of the RFBSDE (96).
For the reflected scheme, the next-step value used in the conditional expectation for the $Z$-component is the post-projection value. Therefore, for $i=N-1,\ldots,0$, we set
[…]
@@
Now, we investigate the errors of the scheme (99) and define the auxiliary quantities as below, for $i=N-1,N-2,\ldots,0$:
@@
$\left\{\begin{aligned} \mathcal{U}_{i}^{m}=&\ \mathbb{E}_{i}\left[\widehat{\mathcal{Y}}_{i+1}^{m}+hf(t_{i},X_{i}^{\pi,m},\mathcal{U}_{i}^{m},\widehat{\mathcal{V}}_{i}^{m})\right],\\ \widehat{\mathcal{U}}_{i}^{m}=&\ \max\left\{\mathcal{U}_{i}^{m},\Phi(X_{i}^{\pi,m})\right\},\\ \widehat{\mathcal{V}}_{i}^{m}=&\ \mathbb{E}_{i}\left[\widehat{\mathcal{Y}}_{i+1}^{m}\frac{(\Delta W_{i}^{m})^{\top}}{h}\right].\end{aligned}\right$. (103)
By the Markov property of the discretized forward process $\{X_{i}^{\pi,m}\}_{0\leq i\leq N}$, we have
$\mathcal{U}_{i}^{m}=\overline{u}(t_{i},X_{i}^{\pi,m}),\qquad\widehat{\mathcal{V}}_{i}^{m}=\widehat{v}(t_{i},X_{i}^{\pi,m}),$ (104)
where $\overline{u}$ and $\widehat{v}$ are some deterministic but unknown functions.
For the Monte Carlo approximation of these conditional expectations, define
@@
$\left\{\begin{aligned} \widetilde{\mathcal{U}}_{i}^{m}:=&\ \frac{1}{K_{i}}\sum_{k=1}^{K_{i}}\widehat{\mathcal{Y}}_{i+1}^{m,k}+hf\left(t_{i},X_{i}^{\pi,m},\widetilde{\mathcal{U}}_{i}^{m},\widetilde{\widehat{\mathcal{V}}}_{i}^{m}\right),\\ \widetilde{\widehat{\mathcal{U}}}_{i}^{m}:=&\ \max\left\{\widetilde{\mathcal{U}}_{i}^{m},\Phi(X_{i}^{\pi,m})\right\},\\ \widetilde{\widehat{\mathcal{V}}}_{i}^{m}:=&\ \frac{1}{K_{i}}\sum_{k=1}^{K_{i}}\widehat{\mathcal{Y}}_{i+1}^{m,k}\frac{(\Delta W_{i}^{m,k})^{\top}}{h}.\end{aligned}\right$. (105)
Let us introduce
@@
$\widehat{\varepsilon}_{i}^{y}=$ $\ \inf_{\widehat{\mathcal{NN}}_{y,i,1+d,n}^{\varrho,L,\mathbf{n}}}\mathbb{E}\left[\left|\overline{u}(t_{i},X_{i}^{\pi,m})-\widehat{\mathcal{NN}}_{y,i,1+d,n}^{\varrho,L,\mathbf{n}}(p_{i}^{m},\theta_{y,i}^{r_{j}})\right|^{2}\right],$ (106)
@@
$\widehat{\varepsilon}_{i}^{z}=$ $\ \inf_{\widehat{\mathcal{NN}}_{z,i,1+d,n\times d}^{\varrho,L,\mathbf{n}}}\mathbb{E}\left[\left|\widehat{v}(t_{i},X_{i}^{\pi,m})-\widehat{\mathcal{NN}}_{z,i,1+d,n\times d}^{\varrho,L,\mathbf{n}}(p_{i}^{m},\theta_{z,i}^{r_{j}})\right|^{2}\right],$
and